# The Kobayashi-Hitchin correspondence for nef and big classes

### 11 Mar 2026 v1

@@ 3.4 Uniqueness and HYM $\Rightarrow$ stability
We will use the following lemma that is shown by the standard computations.
@@
We use the notations in Notation 3.20. Let $h=h_{0}\Psi$ be a $T$-adapted HYM metric on $E$ with $\det\Psi=1$. Assume that $\sup_{X\setminus D}|s_{D}^{m}\Psi|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}^{m}}\leq C$ and
$\int_{X\setminus D}|s_{D}^{m}\partial^{\nabla_{h_{0}}}\Psi|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0},T}^{2}T^{n}\leq C$. Then, for the inverse morphism $\Psi^{-1}$ of $\Psi$, the following holds:
  • $\sup_{X\setminus D}|s_{D}^{m}\Psi^{-1}|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}}\leq C$.
  • $\int_{X\setminus D}|s_{D}^{3m}\partial^{\nabla_{h_{0}}}\Psi^{-1}|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0^{3m}}}^{2}T^{n}\leq C$.
We firstly establish the following lemma.
@@
We use notations in Notation 3.20. Fix $l>0$ sufficently large. Let $h_{D}:=h_{D,0}\Psi_{D}^{2}$ be a $T$-adapted HYM metric on $\mathcal{O}(lD)$ constructed in Theorem 3.17. Let us consider the following composition of holomorphic morphisms:
@@
where $h_{{\rm triv}}$ is the trivial metric on $\mathcal{O}_{X}$ and $\nabla_{D}=\Psi_{D}^{-1}\circ\nabla_{h_{D}}\circ\Psi_{D}$ is the $T$-admissible HYM connection on $(\mathcal{O}(lD),h_{D,0})|_{X\setminus D}$. Then the following holds:
@@
We recall that, by the construction of $\Psi_{D}$ (see Theorem 3.13 and Proposition 3.15), it satisies
@@
Since $s_{D}^{l}:(\mathcal{O}_{X},\bar{\partial}^{\nabla_{h_{{\rm triv}}}})\to(\mathcal{O}(lD),\bar{\partial}^{\nabla_{h_{D,0}}})$ and $\Psi_{D}:(\mathcal{O}(lD),\bar{\partial}^{\nabla_{h_{D,0}}})\to(\mathcal{O}(lD)|_{X\setminus D},\bar{\partial}^{\nabla_{D}})$ is holomorphic (refer to Theorem 3.13) and since $\sqrt{-1}\Lambda_{T}F_{h_{D}}=0$ by Lemma 3.19, we obtain
@@
is a holomorphic isomorphism satisfying, for $m$ in Notation 3.20,
@@
By Theorem 3.17, the holomorphic line bundle $\mathcal{O}((4l+1)D)$ admits a $T$-adapted HYM metric $h_{D}=h_{D,0}\Psi_{D}^{2}$ with $\det\Psi_{D}=1$. Let us define $\nabla_{D}:=\Psi_{D}\circ\nabla_{h_{D}}\circ\Psi_{D}^{-1}$ a $T$-adapted HYM connection of $(\mathcal{O}((4l+1)D),h_{D,0})$. Then, by Theorem 3.13 and Proposition 3.15, we have that $\Psi_{D}$ satisfies
@@
Here $s:=({\rm{id}}_{E}\otimes\Psi_{D})\circ({\rm{id}}_{E}\otimes s_{D}^{4l+1})\circ\Psi_{2}\circ\Psi_{1}^{-1}$ is a holomorphic morphism from $(E,\bar{\partial}^{\nabla_{1}})$ to $(E\otimes\mathcal{O}((4l+1)D),\bar{\partial}^{\nabla_{2}\otimes\nabla_{D}})$. We remark that the $T$-HYM constant of $\nabla_{D}$ equals to 0 by Lemma 3.19. Then the usual calculation shows the following point-wise equation on $X\setminus D$:
@@
We use the notations in Notation 3.20. Assume that a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric $h$. Then $h$ computes the $\alpha^{n-1}$-slope of $E$:
@@
Let $h$ be the $T$-adapted HYM metric on $E$. By Lemma 3.5 and Theorem 3.17, the complex hermitian line bundle $(\det E,\det h_{0})$ admits the $\omega_{i}$-HYM connection $\nabla_{i}$ and the $T$-adapted HYM connection $\nabla_{\det E}$ such that $\nabla_{i}\to\nabla_{\det E}$ locally smoothly on $X\setminus D$. We furthermore have the holomorphic isomorphism $\Psi_{\det E}:(\det E,\bar{\partial}^{E})|_{X\setminus D}\to(\det E|_{X\setminus D},\bar{\partial}^{\nabla_{\det E}})$ appeared in Theorem 3.17. Since $\nabla_{i}$ is $\omega_{i}$-HYM, it satisfies the $\omega_{i}$-HYM equation $\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}}=\lambda_{i}{\rm{id}}$ on $X\setminus D$ with
$\lambda_{i}=\frac{c_{1}(\det E)\cdot\{\omega_{i}\}^{n-1}}{\{\omega_{i}\}^{n}}.$
Since $\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}}\to\sqrt{-1}\Lambda_{T}F_{\nabla_{\det E}}$ at each point in $X\setminus D$, we obtain
$\sqrt{-1}\Lambda_{T}F_{\nabla_{\det E}}=\lambda\hbox{ with $\lambda=\frac{c_{1}(\det E)\cdot\alpha^{n-1}}{\alpha^{n}}.$}$ (3.50)
By Theorem 3.17, the pull-back metric $\Psi_{\det E}^{*}(\det h_{0})$ is a $T$-adapted HYM metric on $\det E$ with the chern connection $\Psi_{\det E}^{-1}\circ\nabla_{\det E}\circ\Psi_{\det E}$. By the assumption, we also have a $T$-adapted HYM metric $\det h$ on $\det E$ induced by the $T$-adapted HYM metric $h$ on $E$. By Theorem 3.23, we have $\Psi_{\det E}^{-1}\circ\nabla_{\det E}\circ\Psi_{\det E}=\nabla_{\det h}$. Then, by (3.50), we obtain
${\rm{Tr}}(\sqrt{-1}\Lambda_{T}F_{h})=\lambda.$
Let $\widetilde{\lambda}$ be the $T$-HYM constant of $h$, that is, it satisfies
$\sqrt{-1}\Lambda_{T}F_{h}=\widetilde{\lambda}{\rm{id}}_{E}.$
Then, by taking trace and integrating over $X\setminus D$ with respect to $T^{n}$, we obtain
${\rm{rk}}E\alpha^{n}\widetilde{\lambda}=\int_{X\setminus D}{\rm{Tr}}(\sqrt{-1}\Lambda_{T}F_{h})T^{n}=\alpha^{n}\lambda.$
Hence, together with (3.50), we obtain the result.
@@
We use the notations in Notation 3.20. Assume that holomorphic vector bundles $E_{1}$ and $E_{2}$ are $\alpha^{n-1}$-slope stable with ${\rm{rk}}E_{1}={\rm{rk}}E_{2}$ and $\mu_{\alpha}(E_{1})=\mu_{\alpha}(E_{2})$. Let $g:E_{1}|_{X\setminus D}\to E_{2}|_{X\setminus D}$ be a nonzero holomorphic morphism satisfying
@@
We use the notations in Notation 3.20. If $E$ is $\alpha^{n-1}$-slope stable, then a $T$-adapted HYM metric on $E$ is unique up to scaling, if it exists.
@@
on $X\setminus D$ since $g$ is positive definite. If we denote the $T$-adapted HYM metric $h_{D}$ on $\mathcal{O}(lD)$ by $h_{D}=h_{D,0}\Psi_{D}^{2}$, then we can assume $|s_{D}\Psi_{D}|_{h_{D,0}\otimes h_{D,0}^{*}}$ is bounded by Theorem 3.13 (2). If we further remark that $h_{1}$ and $h_{2}$ are both $T$-adapted, for sufficiently large $l\in\mathbb{Z}_{>0}$ we have
@@
$\int_{X\setminus D}\left|(\nabla_{h_{1}}\otimes\nabla_{h_{1}}^{*}\otimes\nabla_{h_{D}})\overline{s_{D}^{l}}g\right|_{h_{1}\otimes h_{1}^{*}\otimes h_{D},T}T^{n}<\infty.$ (3.56)
@@
Thus we obtain
@@
Then, by (3.59) and and (3.4), we obtain that
@@
We use the notations in Notation 3.20. If a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric, then $E$ is $\alpha^{n-1}$-slope polystable.
@@
Let $\pi:Y\to X$ be a sequence of blow-ups so that $F:=\pi^{*}\mathcal{F}/{\rm{Tor}}$ is locally free. Then we have $\mu_{\pi^{*}\alpha}(F)=\mu_{\alpha}(\mathcal{F})\geq\mu_{\alpha}(E)=\mu_{\pi^{*}\alpha}(\pi^{*}E)$. We denote by $i:F\to\pi^{*}E$ the induced morphism from the inclusion $\iota:\mathcal{F}\hookrightarrow E$. We remark that $\pi^{*}T$ is an adapted current on $Y$ (Lemma 2.19), and $F$ is $\pi^{*}\alpha^{n-1}$-slope stable (Lemma 2.12), $F$ admits a $T$-adapted HYM metric $h_{F}$. Let $D^{\prime}:=E_{nK}(\pi^{*}\alpha)$ be a snc divisor in $Y$ given by the non-Kähler locus of $\pi^{*}\alpha$. We denote by $h_{D^{\prime}}$ the $\pi^{*}T$-adapted HYM metric on $\mathcal{O}(lD)$ for sufficiently large $l\in\mathbb{Z}_{>0}$. Then we consider the following composition of morphisms of $\pi^{*}T$-adapted HYM vector bundles:
@@
where $\lambda_{E}$ is the $T$-HYM constant of $E$ and $\lambda_{F}$ is the $\pi^{*}T$-HYM constant of $F$. By Corollary 3.24 and (3.60), we have $\lambda_{E}\leq\lambda_{F}$. Since $\pi^{*}h_{E}$, $h_{F}$ and $h_{D^{\prime}}$ are all $\pi^{*}T$-adapted, each term in (3.4) is integrable over $Y\setminus D^{\prime}$ with respect to $T^{n}$ (refer to (3.42)). Furthermore, by the cut-off argument as (3.37), we obtain that
@@
We use the notations in Notation 3.20. If a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric, then it is unique up to scaling.
@@ 3.5 Asymptotic behavior of adapted HYM metrics toward $E_{nK}(\alpha)$
The proof of the estimates of $T$-adapted HYM metrics Theorem 3.13 and the uniqueness of $T$-adapted HYM metric Theorem 3.26, we obtain the following estimates.

### 18 Mar 2026 v2

@@ 3.4 Uniqueness and HYM $\Rightarrow$ stability
In this section, the following lemma will repeatedly used.
@@
We use the notations in Notation 3.18. Let $(F,h_{F})$ be a holomorphic hermitian vector bundle on $X$. Assume that $F$ admits a $T$-adapted HYM metric $h=h_{F}\Psi^{2}$. Let us define a connection $\nabla$ of $F$ by $\nabla=\Psi\circ\nabla_{h}\circ\Psi^{-1}$. Then $\nabla$ is a $T$-admissible HYM connection of a complex hermitian vector bundle $(F,h_{F})$. Furthermore, we have that the morphism $\Psi:(F,\bar{\partial}^{F})\to(F,\bar{\partial}^{\nabla})$ is holomorphic.
Proof.
We firstly prove that $\nabla$ is a $h_{0}$-connection. It is a consequence of the direct computation as follows: For any local section $u$ and $v$ of $F$,
$d\big(h_{F}(u,v)\big)$ $=d\big(h(\Psi^{-1}u,\Psi^{-1}v)\big)$
$=h\big(\nabla_{h}(\Psi^{-1}u),\Psi^{-1}v\big)+h\big(\Psi^{-1}u,\nabla_{h}(\Psi^{-1}v)\big)$
$=h_{F}\big((\Psi\circ\nabla_{h}\circ\Psi^{-1})u,v\big)+h_{F}\big(u,(\Psi\circ\nabla_{h}\circ\Psi^{-1})v\big)$
$=h_{F}(\nabla u,v)+h_{F}(u,\nabla v).$
Next we see that $\Psi:(F,\bar{\partial}^{F})\to(F,\bar{\partial}^{\nabla})$ is holomorphic. Since $\bar{\partial}^{\nabla_{h}}=\bar{\partial}^{F}$, the following computation works by the definition of $\nabla$:
$(\bar{\partial}^{\nabla}\otimes\bar{\partial}^{F^{*}})\Psi=\bar{\partial}^{\nabla}\circ\Psi-\Psi\circ\bar{\partial}^{F}=\Psi\circ\bar{\partial}^{\nabla_{h}}-\Psi\circ\bar{\partial}^{F}=0.$
We recall that for any endomorphism $A$ of $E$, we have at each point
$A^{*_{h}}=h^{-1}\overline{A}^{t}h=\Psi^{-2}(h_{0}^{-1}\overline{A}^{t}h_{0})\Psi^{2}=\Psi^{-2}A^{*_{h_{0}}}\Psi^{2}.$
Hence
$|A|_{h}^{2}={\rm{Tr}}(A\cdot A^{*_{h}})={\rm{Tr}}((\Psi A\Psi^{-1})(\Psi^{-1}A^{*_{h_{0}}}\Psi)).$
Then, since $F_{\nabla}=\Psi\circ F_{h}\circ\Psi^{-1}$, we have that
$|F_{\nabla}|_{h_{0}}^{2}={\rm{Tr}}((\Psi\circ\sqrt{-1}F_{h}\circ\Psi^{-1})(\Psi^{-1}\circ\sqrt{-1}F_{h}^{*_{h_{0}}}\circ\Psi))=|F_{h}|_{h}^{2}.$
Hence, since $h$ is $T$-adapted, we obtain that $\nabla$ is a $T$-admissible HYM connection of $(E,h_{0})$. ∎
Now, we recall the estimates for $T$-adapted HYM metrics of $\alpha^{n-1}$-slope stable vector bundles in Theorem 3.17 (1) and (2) which will be used in this section. We first establish the following lemma.
@@
We use notations in Notation 3.18. Fix $l>0$ sufficently large. Let $h_{D}:=h_{D,0}\Psi_{D}^{2}$ be a $T$-adapted HYM metric on $\mathcal{O}(lD)$ constructed in Theorem 3.17. Let us consider the following composition of holomorphic morphisms:
@@
where $h_{{\rm triv}}$ is the trivial metric on $\mathcal{O}_{X}$ and $\nabla_{D}=\Psi_{D}\circ\nabla_{h_{D}}\circ\Psi_{D}^{-1}$ is the $T$-admissible HYM connection on $(\mathcal{O}(lD),h_{D,0})|_{X\setminus D}$. Then the following holds:
@@
We recall that, by the construction of $\Psi_{D}$ (see Theorem 3.17) , it satisfies
@@
Since $s_{D}^{l}:(\mathcal{O}_{X},\bar{\partial}^{\nabla_{h_{{\rm triv}}}})\to(\mathcal{O}(lD),\bar{\partial}^{\nabla_{h_{D,0}}})$ and $\Psi_{D}:(\mathcal{O}(lD),\bar{\partial}^{\nabla_{h_{D,0}}})\to(\mathcal{O}(lD)|_{X\setminus D},\bar{\partial}^{\nabla_{D}})$ is holomorphic (refer to Lemma 3.19) and since $\sqrt{-1}\Lambda_{T}F_{h_{D}}=0$ by Theorem 3.17, we obtain
@@
is a holomorphic isomorphism (see Lemma 3.19) Furthermore, by definition of $T$-adapted HYM metrics, we have that for $m$ in Notation 3.18,
@@
By Theorem 3.17, the holomorphic line bundle $\mathcal{O}((4l+1)D)$ admits a $T$-adapted HYM metric $h_{D}=h_{D,0}\Psi_{D}^{2}$ with $\det\Psi_{D}=1$. Let us define $\nabla_{D}:=\Psi_{D}\circ\nabla_{h_{D}}\circ\Psi_{D}^{-1}$ a $T$-admissible HYM connection of $(\mathcal{O}((4l+1)D),h_{D,0})$. Then, by Theorem 3.17, we have that $\Psi_{D}$ satisfies
@@
Here $s:=({\rm{id}}_{E}\otimes\Psi_{D})\circ({\rm{id}}_{E}\otimes s_{D}^{4l+1})\circ\Psi_{2}\circ\Psi_{1}^{-1}$ is a holomorphic morphism from $(E,\bar{\partial}^{\nabla_{1}})$ to $(E\otimes\mathcal{O}((4l+1)D),\bar{\partial}^{\nabla_{2}\otimes\nabla_{D}})$. We remark that the $T$-HYM constant of $\nabla_{D}$ equals to 0 by Theorem 3.17. Then the usual calculation shows the following point-wise equation on $X\setminus D$:
@@
We use the notations in Notation 3.18. Assume that a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric $h$. Then $h$ computes the $\alpha^{n-1}$-slope of $E$:
@@
Let $h$ be the $T$-adapted HYM metric on $E$. By Theorem 3.17, the holomorphic line bundle $\det E$ admits a $T$-adapted HYM metric $h^{\prime}{}$ whose $T$-HYM constant $\lambda^{\prime}{}$ satisfies
$\lambda^{\prime}{}=\frac{1}{\alpha^{n}}c_{1}(E)\cdot\alpha^{n-1}.$
Let $h$ be the given $T$-adapted HYM metric on $E$ with $T$-HYM constant $\lambda$. Then it induces a $T$-adapted HYM metric $\det h$ on $\det E$ with $T$-HYM constant ${\rm{rk}}E\lambda$. Then by Theorem 3.21, we have $\nabla_{h^{\prime}{}}=\nabla_{\det h}$, and thus the $T$-HYM constant $\lambda^{\prime}{}$ of $\nabla_{h^{\prime}{}}$ coincides with the $T$-HYM constant ${\rm{rk}}E\lambda$ of $\nabla_{\det h}$. Hence we obtain
$\lambda=\frac{\lambda^{\prime}{}}{{\rm{rk}}E}=\frac{1}{\alpha^{n}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}.$
Then we can prove the first equation of the statement by integrating the $T$-HYM equation $\sqrt{-1}\Lambda_{T}F_{h}=\lambda{\rm{id}}$ over $X\setminus D$ against $T^{n}$. ∎
@@
We use the notations in Notation 3.18. Assume that holomorphic vector bundles $E_{1}$ and $E_{2}$ are $\alpha^{n-1}$-slope stable with ${\rm{rk}}E_{1}={\rm{rk}}E_{2}$ and $\mu_{\alpha}(E_{1})=\mu_{\alpha}(E_{2})$. Let $g:E_{1}|_{X\setminus D}\to E_{2}|_{X\setminus D}$ be a nonzero holomorphic morphism satisfying
@@
We use the notations in Notation 3.18. If $E$ is $\alpha^{n-1}$-slope stable, then a $T$-adapted HYM metric on $E$ is unique up to scaling, if it exists.
@@
on $X\setminus D$ since $g$ is positive definite. If we denote the $T$-adapted HYM metric $h_{D}$ on $\mathcal{O}(lD)$ by $h_{D}=h_{D,0}\Psi_{D}^{2}$, then we can assume $|s_{D}\Psi_{D}|_{h_{D,0}\otimes h_{D,0}^{*}}$ is bounded by Theorem 3.17. If we further remark that $h_{1}$ and $h_{2}$ are both $T$-adapted, for sufficiently large $l\in\mathbb{Z}_{>0}$ we have
@@
$\int_{X\setminus D}\left|(\nabla_{h_{1}}\otimes\nabla_{h_{1}}^{*}\otimes\nabla_{h_{D}})\overline{s_{D}^{l}}g\right|^{2}_{h_{1}\otimes h_{1}^{*}\otimes h_{D},T}T^{n}<\infty.$ (3.57)
@@
Thus, together with the facts that $h_{1}$ is $T$-HYM and the $T$-HYM constant of $h_{D}$ equals to 0 (refer to Theorem 3.17), we obtain
@@
Then, by (3.60) and (3.4), we obtain that
@@
We use the notations in Notation 3.18. If a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric, then $E$ is $\alpha^{n-1}$-slope polystable.
@@
Let $\pi:Y\to X$ be a sequence of blow-ups so that $F:=\pi^{*}\mathcal{F}/{\rm{Tor}}$ is locally free. Then we have $\mu_{\pi^{*}\alpha}(F)=\mu_{\alpha}(\mathcal{F})\geq\mu_{\alpha}(E)=\mu_{\pi^{*}\alpha}(\pi^{*}E)$. We denote by $i:F\to\pi^{*}E$ the induced morphism from the inclusion $\iota:\mathcal{F}\hookrightarrow E$. We remark that $\pi^{*}T$ is an adapted current on $Y$ (Lemma 2.18), and $F$ is $\pi^{*}\alpha^{n-1}$-slope stable (Lemma 2.12), $F$ admits a $T$-adapted HYM metric $h_{F}$ by Theorem 3.17. Let $D^{\prime}:=E_{nK}(\pi^{*}\alpha)$ be a snc divisor in $Y$ given by the non-Kähler locus of $\pi^{*}\alpha$. We denote by $h_{D^{\prime}}$ the $\pi^{*}T$-adapted HYM metric on $\mathcal{O}(lD)$ for sufficiently large $l\in\mathbb{Z}_{>0}$. Then we consider the following composition of morphisms of $\pi^{*}T$-adapted HYM vector bundles:
@@
where $\lambda_{E}$ is the $T$-HYM constant of $h_{E}$ and $\lambda_{F}$ is the $\pi^{*}T$-HYM constant of $h_{F}$. By Corollary 3.22 and (3.61), we have $\lambda_{E}\leq\lambda_{F}$. Since $\pi^{*}h_{E}$, $h_{F}$ and $h_{D^{\prime}}$ are all $\pi^{*}T$-adapted, each term in (3.4) is integrable over $Y\setminus D^{\prime}$ with respect to $T^{n}$ (refer to (3.44)). Furthermore, by the cut-off argument as (3.39), we obtain that
@@
We use the notations in Notation 3.18. If a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric, then it is unique up to scaling.
@@ 3.5 Asymptotic behavior of adapted HYM metrics toward $E_{nK}(\alpha)$
The proof of the estimates of $T$-adapted HYM metrics in Theorem 3.17 and the uniqueness of $T$-adapted HYM metric Theorem 3.24, we obtain the following estimates.